OpenStudy (djbreezy):

Name the y value of the solution to the system of equations. Use substitution: -4x + y = 6 -5x - y = 21

HanAkoSolo (jamierox4ev3r):

you might have to set the two to equal each other -4x+y=6 +4x +4x --------- y=4x+6 then -5x-y=21 +5x +5x --------- -y=5x+21 y=-5x-21

HanAkoSolo (jamierox4ev3r):

then you could do this: -5x-21=4x+6 since each end equals y. Then solve for x. Once you solve for x, plug the x value into one of the equations and solve for y ^^

OpenStudy (djbreezy):

x=−3

HanAkoSolo (jamierox4ev3r):

@DJBreezy does this make sense so far?

HanAkoSolo (jamierox4ev3r):

-5x-21=4x+6 +5x +5x ---------- -21=9x+6 -6 -6 --------- -27=9x /9 /9 ------ x=-3 \(\color{green}\checkmark\) that's right :)

HanAkoSolo (jamierox4ev3r):

now we solve for y -4(-3)+y=6 12+y=6 -12 -12 ------- y=-6

HanAkoSolo (jamierox4ev3r):

you can also check with the other equation -5(-3)-y=21 15-y=21 -15 -15 ------- -y=6 y=-6

HanAkoSolo (jamierox4ev3r):

both equations yield y=-6 therefore, the value of y is -6. Do you understand?

OpenStudy (djbreezy):

so the answer is -6??

OpenStudy (djbreezy):

yes and thx

HanAkoSolo (jamierox4ev3r):

no problem :)